OpenStudy (anonymous):

Find the measure of the acute angle formed by 4x-5y=12 and 6x+11y=20.

OpenStudy (anonymous):

slope of first line= 4/5 second line= -6/11

OpenStudy (anonymous):

\[\tan \theta ={ {m _{1}-m _{2}} \over{1+m _{1}m _{2}}}\]